OpenStudy (anonymous):

the fourth, sixth and eigth terms of an AP add to -6. Find the sixth term

OpenStudy (anonymous):

\[ \cdots 2,-2,-6, \cdots \]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

answer is -2 ?

OpenStudy (anonymous):

ohhhh how did you find it

OpenStudy (anonymous):

please show working thx

OpenStudy (anonymous):

I guessed :D

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

hahaha yes totally you know i think you should change your username to GENIUSFORMATH hehehe

OpenStudy (anonymous):

so a+3d + a +5d + a +7d= -6 3a +15d=-6 divide all by 3 a+ 5d = -2 a +5d = the sixth term

OpenStudy (anonymous):

Great! :)